OpenStudy (anonymous):

Serena Williams made a successful first serve 67% of the time in a wimbledon finals match against her sister venus. If she continues to serve at the same rate the net time they play and serves 6 times in the first game, determine the following probabilities (assume that each serve is independent of each other) a. all 6 serves will be in b. exactly 4 serves will be in c. at least 4 serves will be in Please post work, I already know answers, thanks

OpenStudy (callisto):

Serena's winning rate = 2/3 If she gets all 6 serves win, Probability = (win rate)x(win rate)x(win rate)x(win rate)x(win rate)x(win rate) = (win rate)^6 = (2/3)^6 = 62/729 Hmm, is it correct?